A 5-gallon can of paint costs $15.50 while a 7-gallon can costs $17.00. How much is saved in the purchase of 45 gallons by buying the larger cans?
Answer by richwmiller(17219)

9*15.50=139.50
49 gallons with seven 7 gallon cans
7*17=119
139.50-119=$20.50 savings
buy 47 gallons with six 7 gallon cans + one 5 gallon can
6*17+15.50=117.50
139.50-117.50=$22 savings